Perform the integration by transforming the ellipsoidal region of integration into a spherical region of integration and

then evaluating the transformed integral in spherical coordinates.

(a) RRR

G

x

2 dV , where G is the region enclosed by the ellipsoid 9x

2 + 4y

2 + z

2 = 36.

(b) RRR

G

(y

2 + z

2

) dV , where G is the region enclosed by the ellipsoid x

2

a

2

+

y

2

b

2

+

z

2

c

2

= 1


1
Expert's answer
2021-11-30T14:03:23-0500

"\int\int\int_G x^2dv"


where G is "9x^2+4y^2+z^2=36"


"\frac{x^2}{4}+\frac{y^2}{9}+\frac{z^2}{36}=1"


let "x=2u, y=3v, z=6w"


"(\frac{x}{2})^2+(\frac{y}{3})^2+(\frac{z}{6})^2=1\implies u^2+v^2+w^2=1"


then

"\int\int\int_G x^2dv=\int_u\int_v\int_w f(u,v,w)|j|dudvdw"


where

"f(u,v,w)=4u^2"


"|j|=\frac{\delta(x,y,z)}{\delta(u,v,w)}=\begin{vmatrix}
 x_u & x_v & x_w \\
 y_u & y_v&y_w\\ z_u & z_v& z_w
\end{vmatrix}=\begin{vmatrix}
 2 & 0& 0 \\
 0 & 3&0\\ 0 & 0& 6
\end{vmatrix}=36"



"=\int_u\int_v\int_w 144u^2dudvdw"


let "u=rcos\theta sin\phi, v=rsin\theta sin\phi,w=rcos\theta"

"u^2+v^2+w^2=1\implies r^2=1"


"\int_u\int_v\int_w 144u^2dudvdw=\int_r\int_\theta\int_\phi 144r^2cos^2\theta sin^2\phi(r^2sin\theta)drd\theta d\phi"



"=\int_{r=0}^1\int_{\theta=0}^\pi\int_{\phi=0}^{2\pi} 144r^2(sin\theta cos^2\theta)(sin^2\phi)drd\theta d\phi"


"=144[\frac{r^5}{5}]_0^1[\frac{\phi}{2}-\frac{sin2\phi}{4}]_0^{2\pi}(\int_{\theta=0}^\pi(sin\theta cos^2\theta)d\theta)"




"=144(\frac{1}{5})(\frac{\pi}{2})(\frac{2}{3})=\frac{144\pi}{15}"




"=\frac{144\pi}{15}"